Calculate the initial molar concentration of HF solution with a pH of 2.21.(Kb=1.5e-11)
1
Expert's answer
2021-10-22T03:50:36-0400

2.21= 14-Log(1.5× 10)X


2.21= 14-1.6(X)


1.6X = 14-2.21


1.6X= 11.79


= 7.37mol/L